[4 points) Let xi(t) = {2, x2(t) = t, 23(t) = 1. Orthonormalize x1, 12, x3 in this order in the space L?(-1,1].
Answer #1

produet in de fired by - 1 X₂(t)-{X2, X,) < = t- 2 On L² [1,1], the inner 14,8)=5+ Let, X, (t) = £² X2(+)=+, X: (+) = ] Now C